a lot is in the shape of a right triangle. the hypotenuse is 3m longer than the longer leg. the longer leg is 6m longer than twice the length of the shorter leg. find the lengths of the sides of the lot.
Answer by Cromlix(4381)

Shorter leg = x
Longer leg = 6 + 2x
Hypotenuse = 9 + 2x
(6 + 2x)^2 + x^2 = (9 + 2x)^2
36 + 24x + 4x^2 + x^2 = 81 + 36x + 4x^2
Collect like terms
x^2 - 12x - 45 = 0
(x + 3)(x - 15) = 0
x = -3 (No answer)
x = 15
Shorter leg = 15m
Longer leg = 36m
Hypotenuse = 39m
Proof:
15^2 + 36^2 = 39^2
1521m = 1521m
hope this helps.
:-)